OpenStudy (amorfide):

a rectangular based water tank, 1.5m by 1.2m. what is it's depth? ( or the formula for the depth)

OpenStudy (anonymous):

You only give 2 dimensions. Usually, when dealing with a volume based problem, you have 3 dimensions. the 1.5x1.2m shows the rectangle. You need a z-direction to get the up-down position.

OpenStudy (amorfide):

well i was given DV/DT

OpenStudy (anonymous):

How much water can this tank hold?

OpenStudy (amorfide):

make up an amount, i just need to know how to work out the depth

OpenStudy (anonymous):

V=bhd b=base h=height D=depth \[\frac{dV}{dt}=\frac{dV}{db}\frac{db}{dt}+\frac{dV}{dh}\frac{dh}{dt}+\frac{dV}{dD}\frac{dD}{dt}\] You're missing a lot of information. It's very difficult to understand your problem. To find the depth, you need the volume. If you're given dV/dt, and that value is not equal to 0, that means the volume is changing.

OpenStudy (anonymous):

if your base and height are held constant, then your depth is changing. the equation becomes \[\frac{dV}{dt}=\frac{dV}{db}*0+\frac{dV}{dh}*0+\frac{dV}{dD}\frac{dD}{dt}\]

OpenStudy (amorfide):

thank you very much, great help!